# O0.45cmO0.1cm =

### v1

@@ 7 Strong Packing Dimension Zero
${f^{A}:\mathbb{N}\to\{0,1\}^{*}}$ is a frequent $\mathrm{\Delta}$-$A$ covering of $X$ if $f\in\mathrm{\Delta}$ with oracle $A:\mathbb{N}\to\{0,1\}^{*}$ satisfies the following.
@@
We say that $X$ is $\mathrm{\Delta}$-frequently coverable if there is a functional $F\in\mathrm{\Delta}$ such that $F(A)$ is a frequent $\mathrm{\Delta}$-$A$ covering of $X$ for all infinite $A\subseteq\mathbb{N}$.
@@
4:  Run $F^{O}$ and for each $w$ with $|w|=f_{A}(n)$ where $F^{O}(w)>1$ is found set $c=c-1$,
@@
[…]
@@
$f_{O}(n)=\min\{m\in\mathbb{N}|\prod_{i=0}^{m-1}O(w[0\dots i])\geq 2(n+1)\cdot g_{O}(f_{O}(n-1))\text{ for all }w\in\{0,1\}^{m}\}$
@@
At the computable level, $d$ can be computed to arbitrary precision by computing $d_{i}$ exactly after $i$ is large enough. At the algorithmic level, an algorithm can keep track of the seen outputs and perform the necessary changes from last step when connections are found across different levels.
@@
so $F(O)=d$ is an $O$-supermartingale. Note that there are at most $i+1$ extensions of a $x$ in $A_{i}$ to a $w$ in $A_{i+1}$. Now suppose that $S\in X$ and let $n$ be such that $A_{i}$ contains a prefix of $S$ for all $i\geq n$. Let $r>0$ be such that $d(S[0\dots f_{O}(n)])=r$. Then by construction, for all $k\geq f_{O}(n+i)$ we have
@@
We will prove the contrapositive. Assume no such $M$ exists and let $N$ be any oracle Turing machine. We will construct an $A\subseteq\mathbb{N}$ that $N^{A}$ does not produce a frequent cover.
We do this in stages. At stage $s$, let $A^{s-1}$ be the finite set of naturals in $A$ after stage $s-1$, initially empty, and let $m_{s-1}=\max\{A^{s-1}\}$. Let $n_{s}$ be the amount of naturals needed to finish the next $A_{i}$ along with $A_{i-1}$ if necessary as defined in the third condition of Definition Definition. Now let $m\in\mathbb{N}>m_{s-1}$ be such that the last condition in the following construction holds:
@@
By Lemma 7.2, any $S$ with algorithmic strong packing dimension zero has $K(S[0\dots n]|n)\leq c$ for some $c$ and all $n$. Therefore $S$ is decidable, see for instance [15]. ∎
@@
$w\in f(n)\Rightarrow\exists u\in f(n-1)\text{ with }u\sqsubset w$
@@
$\{S\in\mathbf{C}|\forall n\in\mathbb{N}\enspace\exists w\in\delta(n)\enspace w\sqsubseteq S\}.$
@@
$H_{i}=\{n\in\mathbb{N}|M_{i}(n)\text{ halts within }n^{2}\text{ steps}\}.$
Now remove every other element of $H$ (the odd indices) to get the set
@@
$L_{i}=\cup_{i}\{n_{i},n_{i}+1|n_{i+1}\text{ is odd}\}\bigcup\cup_{i}\{n_{i}|n_{i+1}\text{ is odd}\}.$

### v2

@@ 7 Strong Packing Dimension Zero
$f^{A}:\mathbb{N}\to\{0,1\}^{*}$} is a frequent $\mathrm{\Delta}$-AA covering of $X$ if $f\in\mathrm{\Delta}$ with oracle $A:\mathbb{N}\to\{0,1\}^{*}$ satisfies the following.
@@
We say that $X$ is $\mathrm{\Delta}$-frequently coverable if there is a functional $F\in\mathrm{\Delta}$ such that $F(A)$ is a frequent $\mathrm{\Delta}$-AA covering of $X$ for all infinite $A\subseteq\mathbb{N}$.
@@
4: Run $F^{O}$ and for each $w$ with $|w|=f_{A}(n)$ where $F^{O}(w)>1$ is found set c=c1c=c-1,
@@
in order for the supermartingale condition (3.1) to be satisfied. Therefore each of these $n+1$ strings can be output for the set $A_{i}$. Thus, for every $S\in X$, $F^{O}(S[0...n])>1$ for all but finitely many $n\in\mathrm{Range}(f_{A})$ so it must have a prefix in all but finitely many $A_{i}$.
@@
$f_{O}(n)=\min\{m\in\mathbb{N}\mid\prod_{i=0}^{m-1}O(w[0\dots i])\geq 2(n+1)\cdot g_{O}(f_{O}(n-1))\text{ for all }w\in\{0,1\}^{m}\}$
@@
Note that there are at most $i+1$ extensions of a $x$ in $A_{i}$ to a $w$ in $A_{i+1}$ so this results in a $O$-supermartingale. At the computable level, $d$ can be computed to arbitrary precision by computing $d_{i}$ exactly after $i$ is large enough. At the algorithmic level, an algorithm can keep track of the seen outputs and perform the necessary changes from last step when connections are found across different levels.
@@
so $F(O)=d$ is an $O$-supermartingale. Now suppose that $S\in X$ and let $n$ be such that $A_{i}$ contains a prefix of $S$ for all $i\geq n$. Let $r>0$ be such that $d(S[0\dots f_{O}(n)])=r$. Then by construction, for all $k\geq f_{O}(n+i)$ we have
@@
We will prove the contrapositive. Assume no such $M$ exists and let $N$ be any oracle Turing machine. We will construct an $A\subseteq\mathbb{N}$ such that $N^{A}$ does not produce a frequent cover.
We do this in stages. At stage $s$, let $A^{s-1}$ be the finite set of naturals in $A$ after stage s1s-1, initially empty, and let $m_{s-1}=\max\{A^{s-1}\}$. Let $n_{s}$ be the amount of naturals needed to finish the next $A_{i}$ along with $A_{i-1}$ if necessary as defined in the third condition of Definition Definition. Now let $m\in\mathbb{N}>m_{s-1}$ be such that the last condition in the following construction holds:
@@
By Lemma 7.2, any $S$ with algorithmic strong packing dimension zero has $\mathrm{K}(S[0\dots n]|n)\leq c$ for some $c$ and all $n$. Therefore $S$ is decidable, see for instance [17]. ∎
@@
$w\in\delta(n)\Rightarrow\exists u\in\delta(n-1)\text{ with }u\sqsubset w$
@@
$\{S\in\mathbf{C}\mid\forall n\in\mathbb{N}\kern 5.0pt\exists w\in\delta(n)\kern 5.0ptw\sqsubseteq S\}.$
@@
$H_{i}=\{n\in\mathbb{N}\mid M_{i}(n)\text{ halts within }n^{2}\text{ steps}\}.$
Now remove every other element of $H_{i}$ (the odd indices) to get the set
@@
$L_{i}=\cup_{i}\{n_{i},n_{i}+1\mid n_{i+1}\text{ is odd}\}\bigcup\cup_{i}\{n_{i}\mid n_{i+1}\text{ is odd}\}.$
@@
Remark 7.7.
Franklin et al. [8] defined a sequence $S$ to be anti-complex if for every computable order (non-decreasing and unbounded) $f:\mathbb{N}\to\mathbb{N}$, it is the case that $\mathrm{K}(S[0\dots f(n)])\leq n$ for almost every $n$. Hölzl and Porter [14] showed this is equivalent to $\mathrm{KM}(S[0\dots n])\leq f(n)$ for almost every $n$. In a version of our strong algorithmic packing dimension zero that is changed to only require success with respect to computable odds functions (or equivalently computable prediction orders or computable gauge functions), it is routine to see that this gives an equivalent characterization of anti-complex sequences by an analogue of Lemma 5.4. Similarly, a version of algorithmic strong measure zero only requiring success on computable objects corresponds to i.o. anti-complex [14] sequences where the above characterizations of anti-complex only need to hold for infinitely many $n$.